OpenStudy (anonymous):

Solve the given linear system. 2x-y-z=5 x+y+z=7 3x-2y-3z=1

OpenStudy (anonymous):

do you want to substitute a certain variable for this if not i'll just use z

OpenStudy (anonymous):

i need to find all.. x,y and z

OpenStudy (anonymous):

just for substitution purposes from the 2nd equation you can get z=7-x-y so you can substitute this into the othe two equations to make them two variable equations to then use the elimination method on those two for 1 variable

OpenStudy (anonymous):

so the other two become 2x-y-(7-x-y)=5/2x-y-7+x+=5/3x-7=5/3x=12 3x-2y-3(7-x-y)=1/3x-2y-21+3x+3y=1/6x+y-21=1/6x+y=22

OpenStudy (anonymous):

luckily the first only already has the y removed so you can solve for x which is x=4

OpenStudy (anonymous):

substitute x into the 2nd equation and solve 6(4)+y=22 24+y=22 y=-2

OpenStudy (anonymous):

then plug x and y into the final equation x+y+z=7 4-2+z=7 z+2=7 z=5

OpenStudy (anonymous):

so x=4 y=-2 and z=5

OpenStudy (anonymous):

woah... x+y+z=7 4-2+z=7 z+2=7 <-- why did the 2 become positive? z=5

OpenStudy (anonymous):

4-2=2

OpenStudy (anonymous):

it could have also been written as 2+x=7

OpenStudy (anonymous):

ohh.. i see..

OpenStudy (anonymous):

sorry 2+z=7

OpenStudy (anonymous):

2x-y-(7-x-y)=5/2x-y-7+x+=5/3x-7=5/3x=12 3x-2y-3(7-x-y)=1/3x-2y-21+3x+3y=1/6x+y-21=1/6x+y=22 i dont understand this part

OpenStudy (anonymous):

i was just showing how to simplify it and seperated it by/

OpenStudy (anonymous):

i probably should have explained that sorry i didnt want to have to go line by line too take up so much space

OpenStudy (anonymous):

oh.. its ok.. :)

OpenStudy (anonymous):

still sorry for the confusion do you get it now?

OpenStudy (anonymous):

uhm..

OpenStudy (anonymous):

i remember having a tough time with system equations so don't worry it gets easier

OpenStudy (anonymous):

can you show me how did you find z=7.

OpenStudy (anonymous):

i plugged in the variables found earlier into the middle equation where we got z in the beginning

OpenStudy (anonymous):

2x-y-z=5 x+y+z=7 <----- this one 3x-2y-3z=1

OpenStudy (anonymous):

you could have also plugged them into the expression we used to replace z z=7-x-y and still get the same answer z=7-4-(-2) z=7-4+2 z=5

OpenStudy (anonymous):

oohh.. so z could be oth 7 &5?

OpenStudy (anonymous):

im sorry what are you asking?

OpenStudy (anonymous):

z can be z=7 or z= 5?

OpenStudy (anonymous):

z is 5 the 7 just came from the equation

OpenStudy (anonymous):

you can double check by inserting all the variables into to the equations and see if they agree with what the equations say

OpenStudy (anonymous):

if you used 7 it wouldn't have worked

OpenStudy (anonymous):

hmm..

OpenStudy (anonymous):

do you sort of understand the concept?

OpenStudy (anonymous):

kinda.. because im really slow in english..

OpenStudy (anonymous):

ah that okay lucky math is only letters

OpenStudy (anonymous):

do you know that while you were explaining, i was imagining you were some sort of anime xDD

OpenStudy (anonymous):

lolz not really but I do in fact like anime and manga

OpenStudy (anonymous):

:3 SAO

OpenStudy (anonymous):

hahaha yes

OpenStudy (anonymous):

uhm.. could you answer this? 5x+10y=70 5x+25z=270 10y+25z=300.

OpenStudy (anonymous):

lets get 10y from the first equation so 10y=70-5x you dont need to divide by 10 since the 3rd equation has 10y so you can just directly substitute like this 70-5x+25z=300 -5x+25x=230

OpenStudy (anonymous):

whoops -5x+25z=230 then use elimination of that equation and the 2nd 5x+25z=270 -5x+25z=230 50z=500 z=10

OpenStudy (anonymous):

then you can plug in z into the original 2nd and 3rd equations to find x and y

OpenStudy (anonymous):

where did 5x go?

OpenStudy (anonymous):

5x+25z=270 -5x+25z=230 <--- the first 5x got subtracted by the 5x here

OpenStudy (anonymous):

so it got cancelled

OpenStudy (anonymous):

for the rest of the variables 5x+25z=270 5x+25(10)=270 5x+250=270 5x=20 x=4 10y+25z=300 10y+25(10)=300 10y+250=300 10y=50 y=5

OpenStudy (anonymous):

so x=4 y=5 and z=10

OpenStudy (anonymous):

what did you eat that you can answer this kind of equation xD

OpenStudy (anonymous):

ironic considering I haven't eaten today

OpenStudy (anonymous):

:O why?

OpenStudy (anonymous):

its 12:30am

OpenStudy (anonymous):

how did you get this -5x+25z=230

OpenStudy (anonymous):

perhaps you go to sleep..

OpenStudy (anonymous):

ets get 10y from the first equation so 10y=70-5x you dont need to divide by 10 since the 3rd equation has 10y so you can just directly substitute like this 70-5x+25z=300 -5x+25x=230

OpenStudy (anonymous):

yeah I probably should... but i'll finish this first

OpenStudy (anonymous):

from the substitution of y

OpenStudy (anonymous):

oh.. yea.. thanks :DD you can go to sleep now.. i'll just go through this.. :) thanks man.

OpenStudy (anonymous):

no problem! good luck!

OpenStudy (anonymous):

are you a man? 'cause im not sure :)

OpenStudy (anonymous):

Actually I am female

OpenStudy (anonymous):

well.. that goes for a sis x)) thanks.. and Im a femae too.. thanks you very much.. arigatou

OpenStudy (anonymous):

oyasumi! :)